r+2y [9] 1. Consider the system of equations: = 5 6y - 32 = -1 + 2z = 0 4.2 a) Rewrite the system as a matrix equation AX = Bd) Use adj(A) to calculate A-! Be sure to clearly state any formulas you may use. e) Solve the system using A-1.
Answer #1

Given; Equation is: system of linear xtay 51 by-3z = -1 4.91 +2Z = 0 AX =B Solution a) where 2 x b G-3 x=y B= -1 1 2 Z 2 2 G2 A31 1 -243-0 133 6 A31 - - - 16-20 - 6 A32=- 1- -61X3-02 A-33 = 6 432 = 3 T + 412 A13 12 Adj(A)= -12 24 21 12 A22 A23 2 M31I adj (4) IAL 12 CY - -IZ -12 8 G 2.4- 12 12 2 3 1 2.4 8 G AX =B X = AB 12 -4 - S 1 12 2 2 3 24 8 D 12x5 +4x16x0 6044 64 12 1